# Capturing Polysemanticity with PRISM: A Multi-Concept Feature Description Framework

### v3

@@ Results
Figure 6 presents the results, sorted by PRISM score (lower values indicate higher polysemanticity). We report the average of these ratings as the Human polysemanticity score. In addition, we compute Human Cosine Similarity using the same embedding model (gte-Qwen2-1.5B-instruct sentence transformer [45]) and method used for computing the PRISM polysemanticity score, with the only change being the use of human-generated descriptions instead of model-generated ones. As expected, features with low PRISM scores receive lower human similarity ratings and lower embedding-based similarity. For example, the feature from GPT-2 Small (layer 5, feature 30319) has a low PRISM score (0.30), and is judged by participants to be semantically diverse (Human polysemanticity score: 0.40). In contrast, features with high PRISM scores, such as those from Gemma Scope (features 10531 and 603), show strong human agreement (Human polysemanticity score: 0.90 and 0.80, respectively) and high Human Cosine Similarity (0.79 and 0.75). These findings provide both qualitative and quantitative support for the PRISM metric, showing that it aligns well with the human interpretation of polysemanticity.

### v4

@@ Results
Figure 6 presents the results, sorted by PRISM score (lower values indicate higher polysemanticity). We report the average of these ratings as the Human polysemanticity score. In addition, we compute Human Cosine Similarity using the same embedding model (gte-Qwen2-1.5B-instruct sentence transformer [li-2023-GeneralTextEmbeddings]) and method used for computing the PRISM polysemanticity score, with the only change being the use of human-generated descriptions instead of model-generated ones. As expected, features with low PRISM scores receive lower human similarity ratings and lower embedding-based similarity. For example, the feature from GPT-2 Small (layer 5, feature 30319) has a low PRISM score (0.30), and is judged by participants to be semantically diverse (Human polysemanticity score: 0.40). In contrast, features with high PRISM scores, such as those from Gemma Scope (features 10531 and 603), show strong human agreement (Human polysemanticity score: 0.90 and 0.80, respectively) and high Human Cosine Similarity (0.79 and 0.75). These findings provide both qualitative and quantitative support for the PRISM metric, showing that it aligns well with the human interpretation of polysemanticity.